The level of volcanic activity over the past 24 hours has remained relatively low.
Between 16:00 yesterday and 16:00 today there were 6 volcano-tectonic earthquakes, 1 hybrid earthquake, 1 long period earthquake and 5 rockfalls.
Work continued today on the re-installation of further broadband seismometers.
The Gages tiltmeter, which was damaged at the end of March, was repaired today and is now producing real-time data again along with the other tiltmeters at Long Ground and at Hermitage.
Theodolite data collected on Sunday have now been processed. The very clear conditions mean that the data were very good quality and the height of the top of the spine has now been revised to 1031m above sea level. The total height of the spine at the summit is thought to be about 50m.
A team measured the temperature of the 21st September 1997 pyroclastic flow deposit near Trants this afternoon with a new thermocouple kindly loaned to MVO by the Cascades Volcano Observatory. Three different sites were targeted and the temperatures at each site, at a depth of 2m, were 283, 357 and 284 degrees centigrade.
Dust levels at all monitoring stations were raised today.
Residents are reminded to keep listening to Radio Montserrat for any further information.